In a survey of 1002 adults, a polling agency​ asked, "When you​ retire, do you think you will have enough money to live comfortably or not. Of the 1002 ​surveyed, 530 stated that they were worried about having enough money to live comfortably in retirement. Construct a 90​% confidence interval for the proportion of adults who are worried about having enough money to live comfortably in retirement.
Answer #1

Z for 90% confidence interval = Z​​​​​​0.05 = 1.645

p = 530/1002 = 0.529

Confidence interval = (p + Z​​​​​​0.05 * sqrt(p * (1 - p) / n)

= (0.529 + 1.645 * sqrt(0.529 * (1 - 0.529) / 1002)

= (0.529 + 0.026)

= (0.503, 0.555)